OpenStudy (anonymous):

For the hypothesis test H0: μ = 20 against H1: μ > 20 with variance unknown and n = 10, approximate the P-value for the test statistic t0 = 2.207.

OpenStudy (anonymous):

Oops, ignore that question... You can use this table to approximate the \(p\) value. https://www.easycalculation.com/statistics/t-distribution-critical-value-table.php

OpenStudy (anonymous):

With degrees of freedom \(n-1=10-1=9\) and \(t=2.207\), you can determine a range of \(0.025<p<0.05\). Not sure if you're looking for a better approximation or not.

OpenStudy (anonymous):

@SithsAndGiggles Thank you so much :D